20) Using fiscal policy to avoid the recession of 2009.

GDP in 2009 was roughly $15,000 billion. You learned in Chapter 1 that GDP fell approximately 3 percentage points in 2009. A 3% decrease in GDP is $450 billion.

  1. If the propensity to consume were 0.7 by how much would government spending have to have increased to prevent a decrease in output?

  2. If the propensity to consume were 0.7 by how much would taxes have to have cut to prevent any decrease in output? Explain 1 and 2.
Answer #1

MPC = 0.7
Government spending multiplier = 1/(1-mpc) =

1/(1-0.7)= 3.33

Change in real gdp = multiplier *change in government spending

450 = 3.33* change in government spending

Government spending has to increase by 450/3.33= 135 billion

Tax multiplier = -mpc/(1-mpc) = -0.7/(1-0.7)= -2.33

Change in real gdp =tax multiplier * change in taxes

450 = -2.33 * change in taxes

Taxes has to decrease by 450/2.33= 192.86 billion